solve the quadratic equation by factoring and applying the property "ab=0 if and only if a=0or b=0
x^2+2kx-8k^2

Answer by stanbon(75887)

solve the quadratic equation by factoring and applying the property "ab=0 if and only if a=0or b=0
x^2+2kx-8k^2
------
Rewrite as:
x^2 + 4kx - 2kx - 8k^2
-------
Factor by grouping:
x(x+4k)-2k(x+4k)
-----
Factor:
(x+4k)(x-2k) = 0
Solve using the zero rule:
x+4k = 0 or x-2k = 0
x = -4k or x = 2k
=============================
Cheers,
Stan H.
==============================